OpenStudy (math22):

Twice Bill's number added to 17 is 7 . What is Bill's number?

hero (hero):

17+2x = 7 2x = -10 x = -5

OpenStudy (anonymous):

Always use a letter for your unknown, then set up your equation. 17 + 2x = 7 Solve for x.

OpenStudy (anonymous):

Twice Bill's number = 2x is added to 17: 17 + 2x is 7: is the same as =7 So, put it all together: 17 + 2x = 7 Then, solve for x.

OpenStudy (anonymous):

Please click on good answer for me if I helped you out! Thanks!

OpenStudy (math22):

I know how to write theequation , I don't Know how to solve it